If an ideal gas is at 12.0 degrees Celsius and a pressure of 3.60x104 Pa, calculate the temperature if the gas is compressed to half the volume at a pressure of 9.70x104 Pa.
Answer #1

Here,

let the temperature of the gas is T2

Using ideal gas equation

P1 * V1/T1 = P2 * V2/T2

3.6 *10^4 * V/(12 + 273) = 9.7 *10^4 * 0.50 * V/T2

solving for T2

T2 = 384 K = 111 degree C

the temperature of the gas is 111 degree C